OpenStudy (anonymous):

What is the distance between (2, -1) and (-1, -5) on the coordinate plane? 7 units 6 units 5 units 4 units

OpenStudy (anonymous):

@daniellecrook this is so easy if i had a graph

OpenStudy (anonymous):

6 units

OpenStudy (anonymous):

it is 6 units

OpenStudy (anonymous):

@dlxhazedxlb how did u find out

OpenStudy (anonymous):

i found it out @ivanapen8

terenzreignz (terenzreignz):

lol you guys... I don't think so :)

OpenStudy (anonymous):

@terenzreignz what say it to my face

OpenStudy (anonymous):

is this u

terenzreignz (terenzreignz):

Oh no, I wasn't addressing you, @ivanapen8 Only the guys that said it was 6 units.. and anyway, it'd be nicer if you show how to get it :) Behold the distance formula... \[\Large d = \sqrt{(x_2-x_1)^2 + (y_2-y_1)^2}\]

OpenStudy (anonymous):

@terenzreignz oh ok : - ) im bad at math

terenzreignz (terenzreignz):

plugging in those values results in \[\Large d = \sqrt{(-1-2)^2 + (-5+1)^2}\]\[\Large d = \sqrt{(-3)^2 + (-4)^2}\]\[\Large d = \sqrt{9 + 16}\]\[\Large d = \sqrt{25}\]\[\huge \color{blue}{d=5}\]